# Causal Structure Learning in Hawkes Processes with Complex Latent Confounder Networks

### v1

@@ 3.2 Structure Discovery Through Rank Constraints
In this section, we present theorems that connect statistical properties of discretized Hawkes process data to the structure of the window causal graph, ultimately enabling the identification of the summary causal graph—even in the presence of latent subprocesses. Specifically, we show that under the linear causal model representation with uncorrelated white noise (Eq. 2), the causal structure induces characteristic low-rank patterns in the cross-covariance matrices of observed variables.
To illustrate this, consider the example in Fig. 1. Although the summary causal graph contains directed cycles and self-loops, the corresponding window causal graph is a directed acyclic graph (DAG). Furthermore, by construction, events in the future cannot causally influence events in the past, reflecting the inherent temporal directionality known as Granger causality (shojaie2022granger, ). In this example, all three subprocesses are observed. In the summary causal graph, the subprocess set $\mathcal{P}_{\mathcal{G}}\coloneq\{N_{i}\}_{i\in\{1,2\}}$ forms the parent cause set of $N_{1}$. Correspondingly, in the window graph, we consider the observed variable set $\mathbf{N}_{v}\coloneq\{N_{i}^{(j)}\}^{j\in\{n-m,\dots,n\}}_{i\in\{1,2,3\}}$, where $m$ is chosen to be greater than or equal to the number of effective lag variables. Conditioning on the lagged variable set $\mathbf{P}_{v}\coloneq\{N_{i}^{(j)}\}_{{i\in\{1,2\}}}^{j\in\{n-m,\dots,n-1\}}$, we observe that $N_{1}^{(n)}$ is d-separated from the remaining variables $\mathbf{R}_{v}\coloneq\mathbf{N}_{v}\backslash(\mathbf{P}_{v}\cup N_{1}^{(n)})$. Consequently, the rank of the cross-covariance matrix between $N_{1}^{(n)}\cup\mathbf{P}_{v}$ and $\mathbf{R}_{v}\cup\mathbf{P}_{v}$ is expected to be equal to the size of $\mathbf{P}_{v}$. In the following theorems, we formalize these intuitions and establish a precise connection between rank constraints and the causal structure in the window causal graph, which in turn enables the identification of the summary causal graph, even in the presence of latent subprocesses.
Lemma 3.2 (D-separation and Rank Constraints in the Window Causal Graph).
@@
Proposition 3.3 (Identifying Observed Parent Cause Set).
Consider a PO-MHP consisting of observed subprocesses $\mathcal{O}_{\mathcal{G}}\coloneq\{O_{i}\}_{i=1}^{p}$ and potentially latent subprocesses. For any subprocess $O_{1}$ that is not a single-node subprocess isolated from all others and itself, followings are equivalent:
  • In the summary graph, the set $\mathcal{P}_{\mathcal{G}}\subseteq\mathcal{O}_{\mathcal{G}}$ is the parent cause set of the subprocess $O_{1}$.
@@
Figure 2: Examples of complex causal graphs involving latent confounder subprocesses. (a) A summary causal graph where $O_{1},O_{2},O_{3}$, and $O_{4}$ are observed subprocesses, and $L_{1}$ is a latent confounder subprocess. Self-loops are present on $L_{1},O_{3}$, and $O_{4}$. (b) The corresponding window causal graph for (a), illustrating the discretized causal mechanisms among $O_{1},O_{2}$, and $L_{1}$, considering two effective lag variables. (c) A more complex summary graph where the latent confounder $L_{1}$ connects $O_{1}$ and $O_{2}$ through intermediate latent subprocesses $L_{2}$ and $L_{3}$, respectively. All subprocesses have self-loops, except the intermediate latent subprocesses $L_{2}$ and $L_{3}$. (d) An even more intricate summary graph, compared to (c), where the latent confounder $L_{1}$ connects $O_{1}$ and $O_{2}$ via more complex intermediate latent subprocess paths, and $O_{2}$ also has a directed edge to $L_{1}$. All subprocesses, except for the intermediate latent subprocesses, have self-loops.
The condition in Proposition˜3.3 depends solely on the observed variables and is unaffected by the presence of latent subprocesses. Accordingly, the parent cause set $\mathcal{P}_{\mathcal{G}}$ for any subprocess $O_{1}$ that satisfies the criterion can be uniquely identified. Consequently, we can assert that $O_{1}$ is locally independent of all other subprocesses, conditioning on $\mathcal{P}_{\mathcal{G}}$, regardless of latent subprocesses. Moreover, it is worth noting that the four statements in Proposition˜3.3 are, in fact, valid for any general subprocess $N_{i}$, whether observed or latent. However, this observation has little practical relevance, as the data associated with latent subprocesses are inherently unobservable.
@@
One type of latent subprocess is the intermediate latent subprocess, which may lie on directed paths between an observed subprocess and each of its parent cause set. In general, such intermediate latent subprocesses are unidentifiable, as their effects can be equivalently represented by the observed parent cause set. However, owing to the distinctive causal structure of discretized Hawkes processes, once the observed parent causes of an observed subprocess are identified using Proposition˜3.3, it becomes possible to further quantify the number of intermediate latent subprocesses. Due to space limitations, we defer the details of this result to Appendix˜C.
@@
Another type of latent subprocess is the latent confounder subprocess, which is one of key focuses of our paper. It refers to a latent subprocess that must be included in the parent cause set in order to render its observed effect subprocess locally independent of the others. For example, in Fig. 2(a), we have the structure $O_{1}\leftarrow L_{1}\rightarrow O_{2}$, where $L_{1}$ is a latent confounder of $O_{1}$ and $O_{2}$. In this case, $O_{1}$ and $O_{2}$ are locally independent of each other only when conditioning on $L_{1}$. Proposition˜3.3 cannot identify such latent confounder subprocesses, since their corresponding variables are unobserved. Moreover, it is infeasible to directly infer the number of latent subprocesses from the rank deficiency of the observed cross-covariance matrix, unlike methods for i.i.d. data (Sullivant_2010, ; xie2020generalized, ; huang2022latent, ; jin2023structural, ). This is because, as illustrated in the window causal graph in Fig. 2(b), for each variable $O_{1}^{(n)}$, there are $m$ lagged latent variables $\{L_{1}^{(j)}\}_{j\in{\{n-m,\dots,n-1\}}}$ directly influencing it, where $m$ is unknown and may be as large as several hundred or even thousands in practice. Furthermore, note that Definition˜2.1 does not specify the form of the excitation function. To facilitate the inference of latent confounder subprocesses, we impose the following general constraint on the excitation function.
Condition 1 (Excitation Function).
@@
We consider that the excitation function has the form $\phi_{ij}(s)=a_{ij}w(s),\,\forall i,j\in\{1,\dots,l\}$, where $a_{ij}$ is a constant representing the peer influence between event types $i$ and $j$, and $w(s)$ is a common decay function that depends only on the time lag $s$.
@@
This condition is fairly general. For example, the commonly used exponential decay function $\alpha_{ij}e^{-\beta s}$ (zhou2013learning, ) satisfies this condition. Other examples include normalized linear decay, normalized logistic decay, and similar forms (burt2000decay, ). Additionally, we adopt the rank-faithfulness assumption for the Hawkes process, following standard practice in rank-based methods for i.i.d. data (spirtes2013calculation, ; huang2022latent, ), to rule out degenerate cases of Lebesgue measure zero. Further discussions are provided in Appendix˜D.
@@
Given the excitation function $\phi_{ij}(s)=a_{ij}w(s)$, the excitation coefficients in Eq. 2 are given by: $\theta_{ij}^{(k)}=\int_{(k-1)\Delta}^{k\Delta}\phi_{ij}(s)ds=a_{ij}\int_{(k-1)\Delta}^{k\Delta}w(s)ds$, where the integral term $\int_{(k-1)\Delta}^{k\Delta}w(s)ds$ depends only on the time lag $k$. Consider the summary graph and its corresponding window graph with $m=2$ considered lag variables, as illustrated in Figs. 2(a) and 2(b). According to the linear causal model in Eq. 2, the structural equations for the current variables $O_{1}^{(n)}$ and $O_{2}^{(n)}$ are:
$\begin{bmatrix}O_{1}^{(n)}\\ O_{2}^{(n)}\\ \end{bmatrix}=\mathbf{E}\begin{bmatrix}L_{1}^{(n-1)}\\ L_{1}^{(n-2)}\\ \end{bmatrix}+\begin{bmatrix}\epsilon_{o_{1}}^{(n)}+\theta_{o_{1}}^{(0)}\\ \epsilon_{o_{2}}^{(n)}+\theta_{o_{1}}^{(0)}\\ \end{bmatrix},\hfill\mathbf{E}=\begin{bmatrix}a_{o_{1}l_{1}}\int_{0}^{\Delta}w(s)ds&a_{o_{1}l_{1}}\int_{\Delta}^{2\Delta}w(s)ds\\ a_{o_{2}l_{1}}\int_{0}^{\Delta}w(s)ds&a_{o_{2}l_{1}}\int_{\Delta}^{2\Delta}w(s)ds\\ \end{bmatrix}$. (3)
@@
It is clear that the rank of the coefficient matrix $\mathbf{E}$ is 1. Consequently, the cross-covariance matrix satisfies: $\mathrm{rank}\left(\Sigma_{\{O_{1}^{(n)},O_{2}^{(n)}\},\;\{O_{i}^{(j)}\}_{i\in\{3,4\}}^{j\in\{n-m,\dots,n\}}}\right)=1$. For details, see Proposition˜3.5 and its proof in Appendix˜K. This implies the existence of a single latent confounder subprocess $L_{1}$ acting as the parent cause of both $O_{1}$ and $O_{2}$, such that conditioning on $L_{1}$, $O_{1}$ and $O_{2}$ are locally independent of $O_{3}$ and $O_{4}$.
@@
However, if $O_{1}$ and $O_{2}$ in Figure˜2(a) also have self-loops, the rank of the coefficient matrix is no longer $1$. The self-loops introduce additional indirect causal effects propagated from the lagged latent variables $\{L_{1}^{(j)}\}_{j\in\{n-m,\dots,n-1\}}$ through the observed lagged variables $\{O_{i}^{(j)}\}^{j\in\{n-m,\dots,n-1\}}_{i\in\{1,2\}}$ to the current variables $O_{1}^{(n)}$ and $O_{2}^{(n)}$. Fortunately, since these lagged variables are observed, we can include them in the cross-covariance matrix, resulting in $\mathrm{rank}\left(\Sigma_{\{O_{i}^{(j)}\}_{i\in\{1,2\}}^{j\in\{n-m,\dots,n\}},\;\{O_{i}^{(j)}\}_{i\in\{3,4\}}^{j\in\{n-m,\dots,n\}}\cup\{O_{i}^{(j)}\}_{i\in\{1,2\}}^{j\in\{n-m,\dots,n-1\}}}\right)=2m+1$, where $2m$ corresponds to the observed lagged variables from two observed subprocesses $O_{1}$, $O_{2}$, and $1$ corresponds to the latent confounder subprocess. See Appendix˜E for further details.
@@
Furthermore, in more complex scenarios, as illustrated in Figs. 2(c) and 2(d), the causal pathways from the latent confounder $L_{1}$ to the observed subprocesses $O_{1}$ and $O_{2}$ become increasingly intricate. To handle such settings, we introduce a formalized symmetric path situation, which will serve as a key condition in the subsequent theorems to establish a one-to-one correspondence between the graph structure and its observable statistical properties.
@@
Consider a latent confounder $L_{1}$ and a observed effect subprocess set $\mathcal{O}_{\mathcal{G}1}$. The following conditions define the symmetric path situation:
@@
The structure in Fig. 2(c) satisfies Definition˜3.4, where the latent confounder $L_{1}$ connects $O_{1}$ and $O_{2}$ through the intermediate latent subprocesses $L_{2}$ and $L_{3}$, respectively, both of which do not have self-loops. However, if one intermediate subprocess is removed (e.g., $L_{3}$), the condition in Definition˜3.4 is violated, as the path from $L_{1}$ to $O_{1}$ would then contain one intermediate latent subprocess $L_{2}$, while the path from $L_{1}$ to $O_{2}$ would contain none. Similarly, in the more complex structure shown in Fig. 2(d), the core structure formed by the blue causal edges satisfies Definition˜3.4, and the addition of the green edges still preserves this property. However, adding an extra edge, for instance from $L_{5}$ to $L_{3}$, would break the condition, as it would introduce asymmetry and cycles in the paths. The subsequent theorems will leverage this path condition to formally characterize the graph structure involving the identification of the latent confounder subprocess.
@@
Proposition˜3.5 allows us to infer the existence of a latent confounder from its observed effects. This raises an important question: How can we systematically infer the remaining causal relations involving the inferred latent subprocesses? This challenge is illustrated by the four summary graphs in Fig. 3. In the following, we show how the observed effects can serve as surrogates for their associated latent confounders, enabling the recovery of the remaining causal structure.
Figure 3: Illustrative examples of interactions among inferred latent confounder and the remaining observed subprocesses. In (a)–(c), assume $L_{1}$ has been inferred via its observed effects $\{O_{1},O_{2}\}$. (a) $O_{3}$ causes $L_{1}$. (b) Both $L_{1}$ and $O_{3}$ cause $O_{4}$. (c) $L_{1}$ causes $L_{4}$, where $L_{4}$ can be inferred from $\{O_{3},O_{4}\}$. (d) $L_{1}$ serves as the latent confounder of both latent confounder $L_{2}$ and $L_{3}$.
@@
For each latent subprocess $L_{1}$ inferred from its observed effects $\{O_{1},O_{2}\}$, we define one of its observed effects, denoted as $\mathcal{D}e(L_{1})\coloneq O_{1}$, to serve as an observed surrogate of $L_{1}$. This surrogate is chosen such that there exists a directed path from $L_{1}$ to $\mathcal{D}e(L_{1})$ that does not pass through any other observed subprocesses. We further define $\mathcal{S}ib(\mathcal{D}e(L_{1}))$ as the set of observed siblings of $\mathcal{D}e(L_{1})$, containing all known other observed subprocesses affected by $L_{1}$ through paths that also do not pass through other observed subprocesses.
@@
For any observed subprocess $O_{1}$, we adopt the unified notation $\mathcal{D}e(O_{1})=O_{1}$, and correspondingly, $\mathcal{S}ib(\mathcal{D}e(O_{1}))=\emptyset$. Moreover, $\mathcal{S}ib(\mathcal{D}e(L_{1}))$ represents the minimal set of observed subprocesses required to isolate the local influence of $L_{1}$ on the rest of the system, except through $\mathcal{D}e(L_{1})$.
@@
Theorem 3.7 (Identifying Parent Cause Set with Latent Confounder Involved).

### v2

@@ 4.2 Structure Discovery Through Rank Constraints
In this section, we link second-order statistics of Hawkes data to variables in the window causal graph, which in turn enables identification of the summary causal graph—even with latent subprocesses. Under the linear representation in Eq. 2 with white noise, the causal structure induces characteristic low-rank patterns in cross-covariance matrices of observed discretized variables.
4.2.1 From Observed Parents to Latent Confounder Subprocesses
To illustrate, consider Fig. 1. Although the summary causal graph may contain directed cycles and self-loops, the associated window causal graph is a directed acyclic graph (DAG): by construction, future events cannot causally influence the past, reflecting the intrinsic temporal (Granger-causal) directionality (Shojaie and Fox, 2022). In this example, all three subprocesses are observed. In the summary graph, the parent-cause set of $N_{1}$ is $\mathcal{P}_{\mathcal{G}}\coloneq\{N_{i}\}_{i\in\{1,2\}}$. In the window graph, consider the observed variable set $\mathbf{N}_{v}\coloneq\{N_{i}^{(j)}\}_{i\in\{1,2,3\}}^{j\in\{n-m,\dots,n\}}$, where $m$ is chosen to be at least the number of effective lags. Conditioning on the lagged variable set $\mathbf{P}_{v}\coloneq\{N_{i}^{(j)}\}_{i\in\{1,2\}}^{j\in\{n-m,\dots,n-1\}}$ renders $N_{1}^{(n)}$ d-separated from the remaining variables $\mathbf{R}_{v}\coloneq\mathbf{N}_{v}\setminus(\mathbf{P}_{v}\cup N_{1}^{(n)})$. Consequently, the rank of the cross-covariance between $N_{1}^{(n)}\cup\mathbf{P}_{v}$ and $\mathbf{R}_{v}\cup\mathbf{P}_{v}$ equals $|\mathbf{P}_{v}|$. The theorems below formalize this connection between rank constraints and the window-graph structure, which we then leverage to identify the summary causal graph, including cases with latent subprocesses.
Lemma 4.2 (D-separation and Rank Constraints in the Window Graph).
@@ 4.2.1 From Observed Parents to Latent Confounder Subprocesses
Proposition 4.3 (Identifying Observed Parent-Cause Set).
Consider a PO-MHP with observed subprocesses $\mathcal{O}_{\mathcal{G}}\coloneq\{O_{i}\}_{i=1}^{p}$. The followings are equivalent:
  • In the summary graph, the set $\mathcal{P}_{\mathcal{G}}\subseteq\mathcal{O}_{\mathcal{G}}$ is the parent-cause set of the subprocess $O_{1}$.
@@
Figure 2: Examples of causal graphs with latent confounder subprocesses. (a) Summary graph where $O_{1},O_{2},O_{3},O_{4}$ are observed and $L_{1}$ is latent. (Unlike Figure˜1(d), $O_{1},O_{2}$ are shown without self-loops to simplify the derivation.) (b) Corresponding window causal graph among $O_{1},O_{2}$, and $L_{1}$ with two effective lags. (c) More complex case where $L_{1}$ connects $O_{1},O_{2}$ via intermediate latent subprocesses $L_{2},L_{3}$. All subprocesses have self-loops except $L_{2}$ and $L_{3}$. (d) An even more intricate case, extending (c) with more complex intermediate latent subprocess paths and an additional edge $O_{2}\to L_{1}$. All subprocesses except the intermediate latent ones have self-loops.
The criterion in Proposition˜4.3 involves only observed variables. Whenever an observed subprocess $O_{1}$ satisfies this criterion, its parent-cause set $\mathcal{P}_{\mathcal{G}}$ is uniquely determined (by the minimality conditions in items 2–4), and $O_{1}$ is locally independent of all other observed subprocesses given $\mathcal{P}_{\mathcal{G}}$. As we show next, these four equivalent statements remain valid even with latent subprocesses.
@@
One type of latent subprocess is the intermediate latent subprocess, which, if it exists, may lie on directed paths between an observed subprocess and each of its identified parent-cause set. In general, such intermediates are unidentifiable and typically omitted, since their effects can be absorbed by the identified parent-cause set. Nevertheless, owing to the specific structure of discretized Hawkes processes, once the observed parent causes of an observed subprocess have been identified via Proposition˜4.3, it is possible to quantify the number of intermediate latent subprocesses along these paths. Detailed statements and proofs are deferred to Appendix˜C. Unless otherwise stated, for notational simplicity we do not consider intermediate latent subprocesses into parent-cause set.
@@
Another type of latent subprocess is the latent confounder subprocess, a key focus of this paper. It is a latent subprocess that must be included in the parent-cause set to render its observed effect subprocess locally independent of others. For example, in Fig. 2(a) the structure $O_{1}\leftarrow L_{1}\rightarrow O_{2}$ makes $L_{1}$ a latent confounder of $O_{1}$ and $O_{2}$; in this case, $O_{1}$ and $O_{2}$ are locally independent only when conditioning on $L_{1}$. Proposition˜4.3 will not identify the parent-cause set of either $O_{1}$ or $O_{2}$, since such confounders are unobserved. Furthermore, note that Eq. 1 does not specify the excitation function. We impose the following mild constraint on it to facilitate identification.
Assumption 1 (Excitation Function).
@@
We consider that the excitation function takes the form $\phi_{ij}(s)=a_{ij}w(s),\,\forall i,j\in\{1,\dots,l\}$, where $a_{ij}$ is a constant capturing the peer influence between event types $i$ and $j$, and $w(s)$ is a common decay function depending only on the time lag $s$.
@@
This assumption is quite general. For instance, the widely used exponential decay function $\alpha_{ij}e^{-\beta s}$ (Zhou et al., 2013) falls into this class. Other examples include normalized linear decay, normalized logistic decay, and related forms (Burt, 2000). Moreover, following standard practice in causal discovery (Spirtes, 2013; Huang et al., 2022), we also adopt the rank-faithfulness assumption for Hawkes processes. Intuitively, this assumption rules out pathological parameterizations where causal relationships cannot be identified. It holds generically with infinite data, since the degenerate cases where rank-faithfulness fails constitute a set of Lebesgue measure zero (Spirtes, 2013). Further details are deferred to Appendix˜D.
4.2.2 Characterizing Latent Confounders via Rank Constraints
@@
We now characterize when latent confounder subprocesses leave identifiable low-rank signatures in cross-covariance matrices, moving from illustrative examples to general conditions. Given the excitation function $\phi_{ij}(s)=a_{ij}w(s)$, the excitation coefficients in Eq. 2 are $\theta_{ij}^{(k)}=\int_{(k-1)\Delta}^{k\Delta}\phi_{ij}(s)\,ds=a_{ij}\int_{(k-1)\Delta}^{k\Delta}w(s)\,ds$, where the integral term $\int_{(k-1)\Delta}^{k\Delta}w(s)\,ds$ depends only on the time lag $k$. Consider the summary graph and its corresponding window graph with $m=2$ lagged variables, as illustrated in Figs. 2(a) and 2(b). According to the linear causal model in Eq. 2, the structural equations for the current variables $O_{1}^{(n)}$} and $O_{2}^{(n)}$ are
$\begin{bmatrix}O_{1}^{(n)}\\ O_{2}^{(n)}\\ \end{bmatrix}=\mathbf{E}\begin{bmatrix}L_{1}^{(n-1)}\\ L_{1}^{(n-2)}\\ \end{bmatrix}+\begin{bmatrix}\epsilon_{o_{1}}^{(n)}+\theta_{o_{1}}^{(0)}\\ \epsilon_{o_{2}}^{(n)}+\theta_{o_{2}}^{(0)}\\ \end{bmatrix},\hfill\mathbf{E}=\begin{bmatrix}a_{o_{1}l_{1}}\int_{0}^{\Delta}w(s)\,ds&a_{o_{1}l_{1}}\int_{\Delta}^{2\Delta}w(s)\,ds\\ a_{o_{2}l_{1}}\int_{0}^{\Delta}w(s)\,ds&a_{o_{2}l_{1}}\int_{\Delta}^{2\Delta}w(s)\,ds\\ \end{bmatrix}$. (3)
@@
It is evident that the coefficient matrix $\mathbf{E}$ has rank 1. Consequently, the cross-covariance matrix satisfies $\mathrm{rank}\left(\Sigma_{\{O_{1}^{(n)},O_{2}^{(n)}\},\;\{O_{i}^{(j)}\}_{i\in\{3,4\}}^{j\in\{n-m,\dots,n\}}}\right)=1.$ For details, see Proposition˜4.5 and its proof in Appendix˜K. This indicates a single latent confounder subprocess $L_{1}$ that serves as the parent cause of both $O_{1}$ and $O_{2}$, such that, conditional on $L_{1}$, $\{O_{1},O_{2}\}$ are locally independent of $\{O_{3},O_{4}\}$.
@@
However, if $O_{1}$ and $O_{2}$ in Figure˜2(a) also have self-loops (as illustrated in Figure˜1(d)), the rank of the coefficient matrix is no longer $1$. The self-loops generate additional indirect causal effects propagated from the lagged latent variables $\{L_{1}^{(j)}\}_{j\in\{n-m,\dots,n-1\}}$ through the lagged observed variables $\{O_{i}^{(j)}\}_{i\in\{1,2\}}^{j\in\{n-m,\dots,n-1\}}$ to the current variables $O_{1}^{(n)}$ and $O_{2}^{(n)}$. Since these lagged observed variables are available, we can include them in the cross-covariance matrix, which yields $\mathrm{rank}\!\left(\Sigma_{\{O_{i}^{(j)}\}_{i\in\{1,2\}}^{j\in\{n-m,\dots,n\}},\;\{O_{i}^{(j)}\}_{i\in\{3,4\}}^{j\in\{n-m,\dots,n\}}\cup\{O_{i}^{(j)}\}_{i\in\{1,2\}}^{j\in\{n-m,\dots,n-1\}}}\right)=2m+1,$ where $2m$ corresponds to the lagged variables of the two observed subprocesses $O_{1}$ and $O_{2}$, and $1$ corresponds to the latent confounder subprocess. See Appendix˜E for further details.
@@
Furthermore, in more complex scenarios, as illustrated in Figs. 2(c) and 2(d), the causal pathways from the latent confounder $L_{1}$ to the observed subprocesses $O_{1}$ and $O_{2}$ become increasingly intricate. To address such cases, we formalize the symmetric path situation, which precisely characterizes those graphical configurations that induce rank deficiency in certain sub-covariance matrices of the observed subprocesses. It underpins the subsequent theorems by establishing a one-to-one correspondence between the underlying graph structure and its observable statistical properties.
@@
Consider a latent confounder $L_{1}$ and an observed effect subprocess set $\mathcal{O}_{\mathcal{G}1}$. The following conditions define the symmetric path situation:
@@
The structure in Fig. 2(c) satisfies Definition˜4.4, where the latent confounder $L_{1}$ connects $O_{1}$ and $O_{2}$ through the intermediate latent subprocesses $L_{2}$ and $L_{3}$, respectively, both without self-loops. If one intermediate subprocess is removed (e.g., $L_{3}$), the condition in Definition˜4.4 is violated, since the path from $L_{1}$ to $O_{1}$ would then include one intermediate latent subprocess $L_{2}$, whereas the path from $L_{1}$ to $O_{2}$ would include none. Similarly, in the more complex structure shown in Fig. 2(d), the core structure formed by the blue causal edges satisfies Definition˜4.4, and the addition of the green edges still preserves this property. However, adding an extra edge, for instance, from $L_{5}$ to $L_{3}$, would break the condition, as it would introduce both asymmetry and cycles into the paths. The subsequent theorems will leverage this path condition to formally characterize graph structures involving the identification of latent confounder subprocesses.
@@
Proposition˜4.5 allows us to infer the existence of a latent confounder from its observed effects. This naturally raises the question: How can we systematically infer the remaining causal relations involving these inferred latent subprocesses? This challenge is illustrated in the four summary graphs of Fig. 3. In the following, we show how the observed effects can serve as surrogates for their latent confounders, thereby enabling recovery of the remaining causal structure.
Figure 3: Illustrative examples of interactions among inferred latent confounder and the remaining observed subprocesses. In (a)–(c), assume $L_{1}$ has been inferred from its observed effects $\{O_{1},O_{2}\}$. (a) $O_{3}$ causes $L_{1}$. (b) Both $L_{1}$ and $O_{3}$ cause $O_{4}$. (c) $L_{1}$ causes $L_{4}$, where $L_{4}$ can be inferred from $\{O_{3},O_{4}\}$. (d) $L_{1}$ serves as the latent confounder of both inferred latent confounder $L_{2}$ and $L_{3}$.
@@
For each latent subprocess $L_{1}$ inferred from its observed effects $\{O_{1},O_{2}\}$, we designate one of its observed effects, denoted $\mathcal{D}e(L_{1})\coloneq O_{1}$, as an observed surrogate of $L_{1}$. The surrogate is chosen such that there exists a directed path from $L_{1}$ to $\mathcal{D}e(L_{1})$ that does not traverse any other observed subprocess. We further define $\mathcal{S}ib(\mathcal{D}e(L_{1}))$ as the set of observed siblings of $\mathcal{D}e(L_{1})$, consisting of all known other observed subprocesses influenced by $L_{1}$ through paths that likewise do not pass through any other observed subprocess.
@@
For any observed subprocess $O_{1}$, we adopt the unified notation $\mathcal{D}e(O_{1})=O_{1}$ and, correspondingly, $\mathcal{S}ib(\mathcal{D}e(O_{1}))=\emptyset$. Moreover, $\mathcal{S}ib(\mathcal{D}e(L_{1}))$ captures the minimal set of observed subprocesses required to isolate the local influence of $L_{1}$ on the rest of the system, except through $\mathcal{D}e(L_{1})$.
@@
Theorem 4.7 (Identifying Parent-Cause Set with Latent Confounder Involved).